Find the appropriate form of a particular solution y_{p}(x) using method of undetermined coefficients:

y''+4y'+4y=2xe^{-2x}+8sin(2x)

(do not solve for coefficient constants)
Answer #1

let Solicy + 4y+4y= 2x1-23 +2 31(20) The characteristic equ is 1+49+4=0 = - 4116-411) (4) => A=- 2 - 2 Ye (X) = 4e-2x + Cg